OpenStudy (anonymous):

Here is a word problem i am having trouble with: A train leaves a station heading east at 8 am traveling at a rate of 60 mph. At 9:30 am a 2nd train leaves the same station on a parallel track also heading east traveling at a rate of 900 mph. At what time will the 2nd train catch up to the 1st train

OpenStudy (anonymous):

Distance travelled by first train between 8 and 9 30 = 60 *(3/2) =90m now second train moves with a speed of 840 mph with respect to first train so time taken to travel 90m= 90/840 in minutes it is 90*60/840 find it and add it to 9 30 to get the answer

OpenStudy (anonymous):

I think the speed of the second train is 90 and not 900

OpenStudy (anonymous):

oh ok do you know what time though? its ok if you dont thanks.

OpenStudy (anonymous):

time taken is roughly 6.3 approx so the time is 9 36

OpenStudy (anonymous):

for 90 mph, the answer will be 12.30 pm

OpenStudy (a_clan):

9:36 am